A rose garden is designed by joining a rectangle and a semicircle, as shown below. The rectangle is
31ft long and 23ft wide. If the gardener wants to build a fence around the garden, how many feet of fence are required? (Use the value 3.14 for π, and do not round your answer. Be sure to include the correct unit in your answer.)

Answer by macston(5194)

.
Cannot say exactly, as I cannot see the diagram shown, but in general:
Perimeter of the rectangle -diameter of semicircle + half circumference for semicircle
.
Perimeter of rectangle=2(Length+Width)=2(31ft+23ft)=2(54ft)=108 feet
The length of 1/2 circumference of circle=1/2(pi)(diameter)
Since the diameter of the semicircle is taken away from the perimeter of th rectangle, over all we have:
.
Fence needed=108 feet - diameter of semicircle + 1/2(pi)(diameter of semicircle)